OpenStudy (anonymous):

Jessica is baking a cake. The recipe says that she has to mix 96 grams of sugar to the flour. Jessica knows that 1 cup of this particular sugar has a mass of 128 grams. She added one over two of a cup of sugar to the flour. Should Jessica add more sugar to make the exact recipe or did she go over and by what amount? She went over by two over three of a cup She needs to add two over three of a cup She went over by one over four of a cup She needs to add one over four of a cup

OpenStudy (immings):

She will need to add 1/4 of a cup

OpenStudy (anonymous):

Can you explain it so i can understand

OpenStudy (immings):

96/128 is 3/4 If she added 1/2 she needs to add 1/4 to make it = 3/4

OpenStudy (immings):

so if it is easier to use decimals 3/4 is 0.75 1/2 is 0.5 1/4 is 0.25 0.5+0.25 = 0.75

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Can you help with one more?

OpenStudy (immings):

Sure

OpenStudy (anonymous):

Jacob wants to make a graph to show the relationship between temperatures in Fahrenheit and degrees Celsius. He plots the following points: (11, -11.7), (15, -9.4), (25, -3.9), (35, 1.7), (45, 7.2) He uses the following steps to plot the graph: Label origin as (0, 0) On x-axis label Temperature in Fahrenheit. On y-axis label Temperature in degrees Celsius. The scale on x-axis starts from 0 and goes up to 700 at intervals of 20. The scale on the y-axis starts from -12 and goes up to 12 at intervals of 2. Which of the following best describes the graph? It will not be spread out horizontally across the entire coordinate plane because in Step 4 Jacob selected an incorrect scale on the x-axis. It will not be spread out vertically across the entire coordinate plane because in Step 5 Jacob selected an incorrect scale on the y-axis. It will not be spread out horizontally across the entire coordinate plane because in Step 2 Jacob plotted Fahrenheit instead of Celsius on the axis. It will not be spread out vertically across the entire coordinate plane because in Step 3 Jacob plotted Celsius instead of Fahrenheit on the y-axis.

OpenStudy (immings):

ok so just doing q hick sketch of the graph it looks like A is correct it has an incorrect scale on the x-axis